OpenStudy (anonymous):

The temperature at noon was –12°C. In the next 9 hours, the temperature dropped 8°C. What is the temperature at 9 p.m.?

OpenStudy (anonymous):

-12-8=Y

OpenStudy (anonymous):

This is like.. 2nd grade. Good luck

OpenStudy (anonymous):

sorry i realllyyy can not think today... ive had no sleep and i keep drifting off.

OpenStudy (anonymous):

Ok.. Well like i said.. -12-8= ...

OpenStudy (anonymous):

Ok.. Well like i said.. -12-8= ...

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

wait.. -4